OpenStudy (anonymous):

for triangle ABC side a=18, angle C=90°, and side b=24

OpenStudy (anonymous):

do we have to find the area

OpenStudy (calculusxy):

Do you want to find the length of side C?

OpenStudy (sanjanap):

@calculusxy just give them everything! I guess.

OpenStudy (calculusxy):

@SanjanaP What do you mean?

OpenStudy (sanjanap):

Since they didn't ask a specific one, you could give them the missing side and the angles.

OpenStudy (anonymous):

Wait ignore me.

OpenStudy (anonymous):

i had 30 60 90 in my head lol

OpenStudy (sanjanap):

side c = 30

OpenStudy (calculusxy):

To find the length of the missing side of the triangle you have to do: \[a^{2}+b^{2}=c^{2}\] And then find the square root for c^2.

OpenStudy (sanjanap):

That's what I did. It's 30 24^2 +18^2 = sqrt(900) C=30

OpenStudy (calculusxy):

then that should be correct

OpenStudy (sanjanap):

But there is an equation for finding the other angles. a^2= something bc -sin(Adegrees)